Serena williams serves a tennis ball with a velocity of 35.0 m/s upward. if the ball has a mass of 0.15 kg, what is the kinetic energy of the ball?
1
Expert's answer
2019-02-01T11:06:08-0500

The kinetic energy of the ball

E_k=(mv^2)/2 ,

where m - mass of the ball;

v – velocity of the ball.


E_k=(mv^2)/2= (0.15∙〖35〗^2)/2= 91.875 J


Answer: the kinetic energy of the ball equals 91.875 J.